OpenStudy (anonymous):

how many points of inflection does the function f have on the interval 0≤x≤6 if f"(x)= 2-3sqrt(x)cos^3x ? a)1 b)2 c)3 d)4 e)5

OpenStudy (thomas5267):

Is this the second derivative of the function? \[ f''(x)=2-3\sqrt{x}\cos^3(x) \]

OpenStudy (anonymous):

yes

OpenStudy (thomas5267):

Are you allowed to use graphing calculators?

OpenStudy (anonymous):

no and plus i don't own one :/

OpenStudy (usukidoll):

If we have interval 0 <x< 6 maybe we can create three charts one where we let x <0 like pick x = -1 x = 1 for 0<x<6 and let x = 2pi since it's above 6?

OpenStudy (anonymous):

okay can u just show me so i can make sure i understand?

OpenStudy (thomas5267):

Even Mathematica cannot solve \(f''(x)=0\)... However, it can plot a graph.

ganeshie8 (ganeshie8):

Aren't you given a graph along with the question ?

ganeshie8 (ganeshie8):

@dolphins121

OpenStudy (thomas5267):

This is the graph of the function.

OpenStudy (usukidoll):

I see three of them?

OpenStudy (anonymous):

no there is no graph that comes along with it..@ganeshie8

OpenStudy (usukidoll):

I see an absolute min on the graph

OpenStudy (usukidoll):

and an absolute max

OpenStudy (anonymous):

wait but isn't there a way to solve it without the graph? bcus im not allowed to use a calculator nor am i given a graph..

ganeshie8 (ganeshie8):

Okay maybe your teacher wants you do more work.. sketch the graph or figure it out by some other means. Here I found a different version of this q http://gyazo.com/725b83349d7b30bd49a7d8c2dc379e1a

OpenStudy (usukidoll):

hmmm we need to sketch manually then.. starting with first derivative and then pick an x less than 0, between 0 to 6, and above 6. positive first derivative means increasing negative first derivative is decreasing... and then when you take the second derivative concave up occurs when your result is positive and concave down is when the result is negative. I remembered having to manually draw those graphs.

OpenStudy (usukidoll):

wtheck that graph doesn't have inflection and I know what it looks like @ganeshie8

ganeshie8 (ganeshie8):

i think thats the graph of f''(x).. idk google gave me that haha

ganeshie8 (ganeshie8):

nvm thats the grpah for next question in the problem set :/ looks like we need to graph it manually as you're saying

OpenStudy (thomas5267):

The function provided is the second derivative. The function itself is not integrable by hand either. Mathematica returned this when integrated f''(x) twice. \[ \begin{align*} f(x)&=\frac{1}{48} \left(\sqrt{2 \pi } \left(-81 C\left(\sqrt{\frac{2}{\pi }} \sqrt{x}\right)+54 x S\left(\sqrt{\frac{2}{\pi }} \sqrt{x}\right)+2 \sqrt{3} x S\left(\sqrt{\frac{6}{\pi }} \sqrt{x}\right)\right)\\-\sqrt{6 \pi } C\left(\sqrt{\frac{6}{\pi }} \sqrt{x}\right)+6 \sqrt{x} \left(8 x^{3/2}+27 \cos (x)+\cos (3 x)\right)\right) \end{align*} \]

OpenStudy (usukidoll):

omg. yo teacher is evil! for assigning that shiz

OpenStudy (thomas5267):

C is the Fresnel C and S is the Fresnel S. http://en.wikipedia.org/wiki/Fresnel_integral

OpenStudy (anonymous):

ugh it's just so confusing. & haha yeah idk why she assigned a problem this hard :/ @UsukiDoll

OpenStudy (thomas5267):

I finally got the Mathematica to work for me. x=5.43173 is the x coordinate of the inflection point.

OpenStudy (thomas5267):

I think you could only use the graph. I don't think there is another way. Maybe Newton-Raphson but I am not familiar with it.

OpenStudy (anonymous):

oh. well thanks for trying :) i appreciate it

OpenStudy (perl):

@ganeshie8 what problem set are you looking at "nvm thats the graph for next question in the problem set :/ looks like we need to graph it manually as you're saying "

OpenStudy (perl):

just out of curiosity :)

OpenStudy (perl):

it might be easier to solve when is f ' ' (x) > 0 , and f ' ' (x) < 0 , you can do this approximately

OpenStudy (thomas5267):

How do you solve it approximately? I am interested.

OpenStudy (perl):

I graphed f ' ' (x) using a calculator, and there is a zero between 5 and 6

OpenStudy (perl):

Since f ' ' is continuous on [0,6] and since f ' ' (5) = 1.84688 > 0 f ' ' ( 6) = -4.504916 < 0 By the intermediate value theorem there is some value c between 5 and 6 such that f ' ' (c) = 0. And since the sign changes there, it is an inflection point

OpenStudy (anonymous):

so there is only one inflection point?

OpenStudy (perl):

yes because f ' ' (x) stays positive between 0 and 5

OpenStudy (perl):

actually the 'zero' is x = 5.437318

OpenStudy (anonymous):

alright well thanks for ur help. I appreciate it:)

OpenStudy (perl):

I couldn't find this question online to compare it to.

OpenStudy (perl):

thanks @ganeshie8 if you really want to be absolutely there is only one inflection point, you would have to take the derivative of f ' ' (x) , and see where it is increasing/decreasing. ie, you would you have to find f ' ' ' (x) which i am going to do since i hate loose ends :)

OpenStudy (perl):

I'm still curious I googled the phrase "How many points of inflection does the function f have on the interval 0<= x <= 6" and i didnt get that AP calc book. You have some amazing searching skills :)

OpenStudy (perl):

looks like you searched right into google books